Geometric representation of S={Im(z-i)=|z+I|}


1
Expert's answer
2022-01-19T18:10:26-0500

1. Let "z=x+iy, x, y \in \R"

Then


"Im(z-i)=Im(x+iy-i)=y-1"

"|z+i|=|x+iy+i|=\sqrt{x^2+(y+1)^2}"

We have the equation

"y-1=\sqrt{x^2+(y+1)^2}"

"\sqrt{x^2+(y+1)^2}\geq0=>y\geq1"

If "y\geq1," then

"\sqrt{x^2+(y+1)^2}\geq y+1>y-1"

Therefore, the equation


"y-1=\sqrt{x^2+(y+1)^2}"

has no solution.


2. Let "z=x+iy, x, y \in \R"

Then


"Im(z-i)=Im(x+iy-i)=y-1"

"|z+1|=|x+1+iy|=\sqrt{(x+1)^2+y^2}"

We have the equation

"y-1=\sqrt{(x+1)^2+y^2}"

"\sqrt{(x+1)^2+y^2}\geq0=>y\geq1"

If "y\geq1," then

"\sqrt{(x+1)^2+y^2}\geq y>y-1"

Therefore, the equation


"y-1=\sqrt{(x+1)^2+y^2}"

has no solution.